# Robust Measurement of Stellar Streams Around the Milky Way: Correcting Spatially Variable Observational Selection Effects in Optical Imaging Surveys

### v2

@@ Abstract
facilities: Blanco (DECam)software: astropy (Astropy Collaboration et al., 2013), fitsio [^3] https://github.com/esheldon/fitsio, HEALPix (Górski et al., 2005), healpy (Zonca et al., 2019), healsparse [^4] https://github.com/LSSTDESC/healsparse, matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), skyproj [^5] https://github.com/LSSTDESC/skyproj, ugali (Bechtol et al., 2015; Drlica-Wagner et al., 2020)
@@ 1 Introduction
Stellar streams are the tidally disrupting remnants of star clusters and satellite galaxies (Newberg & Carlin, 2016). The discovery and characterization of these systems around the Milky Way has been enabled by wide-field astronomical surveys, initially through matched-filter searches of photometric data (e.g., SDSS; Odenkirchen et al. 2001, Pan-STARRs; Bernard et al. 2014, Grillmair 2017, DES; Shipp et al. 2018), and subsequently using combinations of photometric, astrometric and spectroscopic data (e.g., Gaia; Malhan et al. 2018, S5; Li et al. 2019). Currently, there are more than 120 identified stellar streams around our Galaxy (Mateu, 2023; Bonaca & Price-Whelan, 2025).
@@
Within an individual stream, the stellar positions and velocities are probes of the local acceleration field experienced by the stream stars as they orbit around the Milky Way’s gravitational potential (Bovy et al., 2017). Therefore, in a Milky Way potential absent of substructure, it is expected that streams would show less substructure; although some will be present due to variations in the stripping rate and the formation of epicyclic overdensities (Küpper et al., 2010). But, the presence of dark subhalos passing by a stream will induce perturbations in the on-sky positions, density, and velocity distribution of stream stars. This can result in the formation of gaps and other small scale deviations in the stream track (e.g., Erkal et al. 2016, Bonaca et al. 2019, Banik et al. 2021a, Banik et al. 2021b, Delos & Schmidt 2022).
@@
The population of kinematically cold stellar streams (i.e., ones with globular cluster progenitors) comprise the best probe of this effect due to their small intrinsic velocity dispersions and stream widths (Lu et al., 2025). One such kinematically cold stellar stream is the Phoenix stream. This substructure, originally discovered in early DES data (Balbinot et al., 2016), has a length of 4.6 kpc ($\sim 15^{\circ}$), a width of $0.14^{\circ}$, and a measured velocity dispersion of $\sigma_{rv}=2.66\,\mathrm{km/s}$ (Shipp et al., 2018; Wan et al., 2020). Interestingly, this stream also shows small scale density fluctuations, making it a promising candidate for probing the low mass regime of the subhalo mass function (Tavangar et al., 2022). Due to these characteristics, we use this stream as an example application in our analysis. For thin streams in general, there has been a large effort to follow up discoveries using deeper imaging surveys to access fainter stars and provide the best constraints on the track and density variations (e.g., Erkal et al., 2017; Koposov et al., 2019; Bonaca et al., 2020; Li et al., 2021; Tavangar et al., 2022; Patrick et al., 2022), while proper motion measurements (Shipp et al., 2019) and targeted spectroscopy of brighter members (Bonaca et al., 2021; Li et al., 2022) can be used to refine the stellar samples, constrain dynamics and further characterize the interaction history of these substructures.
@@
One common approach to quantify the observable effects of gravitational interactions with dark matter is through power-spectrum-type analyses of density variations (stars/deg) along a stream (Banik et al., 2021a; Delos & Schmidt, 2022). Such inferences rely on accurate measurements of the density and track along the stream. At small spatial scales, shot noise from the limited number of stars dominates the stream density power spectrum, thus it is most important to reduce bias on measurements at larger scales (tens of deg; Bovy et al., 2017).
@@
In addition, for most optical surveys, the observations are taken over multiple epochs under different observing conditions. This produces variations in detection rates for both stars and galaxies in imaging data. It also causes variations in how often stars and galaxies are incorrectly classified. Already with DES Year 3 data, we find that spatially variable observational selection effects across the survey footprint induce statistically significant effects (Section 4). As our data and analyses become more sophisticated, percent-level variations in the survey-transfer function can induce systematic errors that are much larger than the statistical errors and therefore limit the power of analyses (Everett et al., 2022).
Traditionally, to avoid these systematic biases, a selection for high signal-to-noise objects is applied to minimize variations in observational selection effects. For example, in the DES DR1 stream search, the stellar selection was limited to $g<23.5$ (Shipp et al., 2018). However, at the magnitudes that will be accessible to future surveys, such as the NSF-DOE Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST, $5\sigma\,g$-band depth $\sim 27.4$ corresponding to around 2.3 magnitudes deeper than DES DR1; Ivezić et al., 2019), galaxies are much more common, and even small variations in galaxy misclassification rates could produce large effects on the stellar sample (Tsiane et al., 2025). Therefore, new tools are needed to fully leverage the potential LSST-like data.
In the context of large-scale structure analyses for cosmology, galaxy weight maps have been derived to account for spatially variable observational selection effects that would otherwise produce systematic errors in galaxy clustering measurements. These weight maps are derived using the ansatz that galaxies are isotropically distributed on large angular scales, and that empirical relations between observed galaxy densities and survey properties (e.g., seeing, integrated exposure time) can be used to learn the observational selection function for galaxies across the survey footprint (see Section 5.3 of Rodríguez-Monroy et al., 2022). This approach is not possible for stellar samples, because the intrinsic distribution of stars across the survey footprint is highly non-isotropic.
@@
The paper is organized as follows: in Section 2 we discuss the data from DES Y3 and Balrog that will be used in our corrections. In Section 3 we discuss the calculations involved in our corrections. In Section 4 we apply our corrections to the full DES Y3 footprint and compare these results to pre-corrected data. In Section 5 we test the overall corrective power of our algorithm, how this power changes with larger training sets, and how repeatable our algorithm is. In Section 6 we apply our corrections to observations of the Phoenix stream and investigate the resulting changes in linear density. We then apply our corrections to simulated stellar streams and compare density power spectra. Finally, we conclude and motivate potential future works in Section 7. All code used in this project is available on github [^1] https://github.com/Kyle-Boone/ssi_corrections_des_y3_balrog.
2 Data
In this section we describe the observations, DES Y3 Gold catalog (Section 2.1), and synthetic source injection runs, Balrog (Section 2.2), used in our analysis.
2.1 DES DR1 & Y3 GOLD
@@
2.2 Balrog Synthetic Sources
To compliment the Y3 GOLD data, the synthetic source catalog generated by Balrog (Everett et al., 2022; Suchyta et al., 2016) is used. Balrog is a software package that synthetically injects sources into individual DES images before coaddition and processes them with the same DESDM pipelines as observations. Balrog is the name used to refer to both the software and synthetic object catalog, but for the remainder of this work we will use this name only to refer to the catalog. The injected sources were chosen to be empirical populations of artificial galaxies and stars taken from the deeper DES Deep Field observations (Hartley et al., 2021). Additionally, delta-function stars (delta functions convolved with a local PSF) are injected, which we use for our synthetic stellar sample to avoid any potential galaxy contamination from incorrect Deep Field classifications. These synthetic objects are subject to the same systematic impacts from survey properties as physical objects would be. The 7.4M Balrog objects used for this work (90% Deep Field objects, 10% delta-stars) were injected on a uniform grid on 2,041 randomly chosen tiles out of the 10,338 total Y3 tiles (for a coverage map see Figure 6 of Everett et al. 2022, and more information on the injected objects can be found in their Section 3).
@@ 2.2 Balrog Synthetic Sources
3 Methods
3.1 Overview and Notation
@@
Our algorithm uses SSI objects and survey property maps to derive a relation between the survey properties and the likelihood that a given object will be detected and classified correctly. This will allow us to take catalog level data ($C_{S}$ and $C_{G}$) and correct it based on the survey properties at a given on-sky location. To get this correction, we first use SSI objects to obtain the probability that a detected SSI star/galaxy will be classified correctly ($P\left(IC_{S}|IO_{S}\right)$ and $P\left(IC_{G}|IO_{G}\right)$). We assume that the probability that a detected star (galaxy) will be classified correctly is the same as that probability for an SSI star (galaxy):
@@
Classification probabilities and relative detection rates are used to obtain final corrected counts at each position ($FC_{S}$ & $FC_{G}$). We present two different types of corrections: one estimates what $C_{S}$ would have been with uniform survey properties, and one estimates what $O_{S}$ would have been with uniform survey properties. More details are given in Section 3.3.
@@
3.2 Probability Calculations
Figure 2: This figure shows Balrog delta star correct classification rates before (blue) and after (orange) corrections as a function of the effective exposure time sum in the $i$-band. The green histogram shows the distribution of effective exposure time in the $i$-band across the DES footprint. Training was performed on an 80% subsample of the Balrog objects with the remaining sources used for testing. As well as magnitude cuts, a color cut to the Phoenix isochrone is made (see Section 4.1). A drop in standard deviation after corrections shows a mitigation of some of the variations caused by this survey property.
@@ 3.2 Probability Calculations
As a specific example to illustrate our method, we present $P\left(IC_{S}|IO_{S}\right)$ in our faintest magnitude bin as a function of a single survey property in Figure 2. The green histogram shows the distribution of effective exposure time in the $i$-band for DES Y3. The blue (orange) line shows the ratio before (after) corrections of correctly classified SSI stars to all detected SSI stars as a function of the survey property value (Table 1). Since we only care about the variations in this rate, we divide each line by its average to center them at $1$. For corrections, this ratio of correctly classified SSI stars to all detected SSI stars is computed for all survey properties. We then take the survey property with the largest variation in the ratio and apply an empirical correction. This process is repeated iteratively until one of our termination criteria is met: either $\sigma<0.01$ for all survey properties or, to avoid overfitting and/or long runtimes, after 150 iterations (see Section 5.3). To ensure that we are not just removing noise, we take 80% of our SSI stars to be a training sample and the remaining 20% to be a testing sample. All corrections were calculated on the training sample but then applied to the testing sample to make the plot. In this way, the drop in standard deviation from $\sigma=0.07$ to $\sigma=0.02$ indicates that our correction pipeline has reduced the impact variations in this survey property have on the stellar classification rate. We see $\sigma>0.01$ after corrections since the test set is independent from the training set.
@@
3.3 Algorithm Design
With classification probabilities and relative detection rates established, we now detail the correction algorithms introduced in Section 3.1, which share identical steps except for their final stage.
The first step uses $P\left(C_{S}|O_{S}\right)$, $P\left(C_{G}|O_{G}\right)$, $C_{S}$, and $C_{G}$ to estimate the number of true stars (galaxies) that were detected, $RO_{S}$ ($RO_{G}$). We use a maximum likelihood approach to estimate $RO_{S}$ and $RO_{G}$, which gives:
@@ 3.3 Algorithm Design
4 Correction of the DES Data
In this section we present the results of applying our correction algorithm to DES data. To select a realistic dataset for this pipeline that will be applicable to a stream, we follow the study of the Phoenix stream by Tavangar et al. (2022). Phoenix is a $15^{\circ}$ long, $0.16^{\circ}$ wide, and dynamically cold stellar stream in the Southern Hemisphere (Balbinot et al., 2016; Shipp et al., 2018; Tavangar et al., 2022). Phoenix shows density fluctuations on small scales, making it a well-suited candidate for studying potential perturbations. In addition to its intrinsic properties, the stream lies in the middle of the DES footprint, near a prominent survey depth feature at a right ascension (RA) of $\sim 30$ deg (Figure 1). This feature has increased depth and is one of the most readily apparent survey property features in the entire DES footprint. Therefore, we can use observations of the Phoenix stream to demonstrate the application of our pipeline and its effect on the generation of density maps.
@@ 4 Correction of the DES Data
4.1 Matched Filter Data selection
To derive corrections for a given stellar or galactic density map we want to place the same observational selection criteria on the injected objects as we would place on observations. In the case of the Phoenix stream, we use the matched filter from Tavangar et al. (2022). This filter is described in detail in their analysis, but generally it uses a synthetic isochrone to generate a selection region in color-magnitude space. The size of this region is defined by the expected width of the stellar population convolved with observational uncertainties. We use an isochrone from Bressan et al. (2012) as implemented in ugali (Bechtol et al., 2015; Drlica-Wagner et al., 2020) with an age of a =12.8=12.8 Gyr, metallcity of [Fe/H] =2.5=-2.5, and distance of d =17.4=17.4 kpc (Tavangar et al., 2022).
We also use a magnitude limit of $r<24.5$, where completeness falls to $\sim 20\%$ (see Everett et al. 2022 Figure 9). This can be compared to Shipp et al. (2018) which use the same observations (DES Y3) but place a limit around one magnitude brighter ($g<23.5$). It is also $0.3$ mag fainter than the $r<24.2$ limit used in Tavangar et al. (2022) which uses DES Y6 data (roughly twice as many observations as DES Y3).
As mentioned in Section 2.1, in our correction algorithm we apply corrections to different magnitude bins. Our magnitude bins are as follows: $r\leq 22.9$, $22.9<r\leq 23.9$, $23.9<r\leq 24.5$. These magnitude bins were chosen as they had similar relative detection rate variations in testing sets when these variations were extrapolated out to expected values for a more current Balrog run. For more details, refer to Appendix E. These selections were placed in addition to the ones described in Section 2.
4.2 Algorithm Adjustments for Real Data

### v3

@@ Abstract
facilities: Blanco (DECam)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), fitsio [^3] https://github.com/esheldon/fitsio, HEALPix (Górski et al., 2005), healpy (Zonca et al., 2019), healsparse [^4] https://github.com/LSSTDESC/healsparse, matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), skyproj [^5] https://github.com/LSSTDESC/skyproj, ugali (Bechtol et al., 2015; Drlica-Wagner et al., 2020)
@@ I Introduction
Stellar streams are the tidally disrupting remnants of star clusters and satellite galaxies (Newberg and Carlin, 2016). The discovery and characterization of these systems around the Milky Way has been enabled by wide-field astronomical surveys, initially through matched-filter searches of photometric data (e.g., SDSS; Odenkirchen et al. 2001, Pan-STARRs; Bernard et al. 2014, Grillmair 2017, DES; Shipp et al. 2018a), and subsequently using combinations of photometric, astrometric and spectroscopic data (e.g., Gaia; Malhan et al. 2018, S5; Li et al. 2019). Currently, there are more than 120 identified stellar streams around our Galaxy (Mateu, 2023; Bonaca and Price-Whelan, 2025).
@@
Within an individual stream, the stellar positions and velocities are probes of the local acceleration field experienced by the stream stars as they orbit around the Milky Way’s gravitational potential (Bovy et al., 2017). Therefore, in a Milky Way potential absent of substructure, it is expected that streams would show less substructure; although some will be present due to variations in the stripping rate and the formation of epicyclic overdensities (Küpper et al., 2010). But, the presence of dark subhalos passing by a stream will induce perturbations in the on-sky positions, density, and velocity distribution of stream stars. This can result in the formation of gaps and other small scale deviations in the stream track (e.g., Erkal et al. 2016, Bonaca et al. 2019, Banik et al. 2021b, Banik et al. 2021c, Delos and Schmidt 2022).
@@
The population of kinematically cold stellar streams (i.e., ones with globular cluster progenitors) comprise the best probe of this effect due to their small intrinsic velocity dispersions and stream widths (Lu et al., 2026). One such kinematically cold stellar stream is the Phoenix stream. This substructure, originally discovered in early DES data (Balbinot et al., 2016), has a length of 4.6 kpc ($\sim 15^{\circ}$), a width of $0.14^{\circ}$, and a measured velocity dispersion of $\sigma_{rv}=2.66\,\mathrm{km/s}$ (Shipp et al., 2018a; Wan et al., 2020). Interestingly, this stream also shows small scale density fluctuations, making it a promising candidate for probing the low mass regime of the subhalo mass function (Tavangar et al., 2022). Due to these characteristics, we use this stream as an example application in our analysis. For thin streams in general, there has been a large effort to follow up discoveries using deeper imaging surveys to access fainter stars and provide the best constraints on the track and density variations (e.g., Erkal et al., 2017; Koposov et al., 2019; Bonaca et al., 2020; Li et al., 2021; Tavangar et al., 2022; Patrick et al., 2022), while proper motion measurements (Shipp et al., 2019) and targeted spectroscopy of brighter members (Bonaca et al., 2021; Li et al., 2022) can be used to refine the stellar samples, constrain dynamics and further characterize the interaction history of these substructures.
@@
One common approach to quantify the observable effects of gravitational interactions with dark matter is through power-spectrum-type analyses of density variations (stars/deg) along a stream (Banik et al., 2021b; Delos and Schmidt, 2022). Such inferences rely on accurate measurements of the density and track along the stream. At small spatial scales, shot noise from the limited number of stars dominates the stream density power spectrum, thus it is most important to reduce bias on measurements at larger scales (tens of deg; Bovy et al., 2017).
@@
In addition, for most optical surveys, the observations are taken over multiple epochs under different observing conditions. This produces variations in detection rates for both stars and galaxies in imaging data. It also causes variations in how often stars and galaxies are incorrectly classified. Already with DES Year 3 data, we find that spatially variable observational selection effects across the survey footprint induce statistically significant effects (Section IV). As our data and analyses become more sophisticated, percent-level variations in the survey-transfer function can induce systematic errors that are much larger than the statistical errors and therefore limit the power of analyses (Everett et al., 2022).
Traditionally, to avoid these systematic biases, a selection for high signal-to-noise objects is applied to minimize variations in observational selection effects. For example, in the DES DR1 stream search, the stellar selection was limited to $g<23.5$ (Shipp et al., 2018a). However, at the magnitudes that will be accessible to future surveys, such as the NSF-DOE Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST, $5\sigma\,g$-band depth $\sim 27.4$ corresponding to around 2.3 magnitudes deeper than DES DR1; Ivezić et al., 2019), galaxies are much more common, and even small variations in galaxy misclassification rates could produce large effects on the stellar sample (Tsiane et al., 2025). Therefore, new tools are needed to fully leverage the potential LSST-like data.
In the context of large-scale structure analyses for cosmology, galaxy weight maps have been derived to account for spatially variable observational selection effects that would otherwise produce systematic errors in galaxy clustering measurements. These weight maps are derived using the ansatz that galaxies are isotropically distributed on large angular scales, and that empirical relations between observed galaxy densities and survey properties (e.g., seeing, integrated exposure time) can be used to learn the observational selection function for galaxies across the survey footprint (see Section 5.3 of Rodríguez-Monroy et al., 2022b). This approach is not possible for stellar samples, because the intrinsic distribution of stars across the survey footprint is highly non-isotropic.
@@
The paper is organized as follows: in Section II we discuss the data from DES Y3 and Balrog that will be used in our corrections. In Section III we discuss the calculations involved in our corrections. In Section IV we apply our corrections to the full DES Y3 footprint and compare these results to pre-corrected data. In Section V we test the overall corrective power of our algorithm, how this power changes with larger training sets, and how repeatable our algorithm is. In Section VI we apply our corrections to observations of the Phoenix stream and investigate the resulting changes in linear density. We then apply our corrections to simulated stellar streams and compare density power spectra. Finally, we conclude and motivate potential future works in Section VII. All code used in this project is available on github [^1] https://github.com/Kyle-Boone/ssi_corrections_des_y3_balrog and Zenodo (Boone, 2026).
II Data
In this section we describe the observations, DES Y3 Gold catalog (Section II.1), and synthetic source injection runs, Balrog (Section II.2), used in our analysis.
II.1 DES DR1 & Y3 GOLD
@@
II.2 Balrog Synthetic Sources
To compliment the Y3 GOLD data, the synthetic source catalog generated by Balrog (Everett et al., 2022; Suchyta et al., 2016) is used. Balrog is a software package that synthetically injects sources into individual DES images before coaddition and processes them with the same DESDM pipelines as observations. Balrog is the name used to refer to both the software and synthetic object catalog, but for the remainder of this work we will use this name only to refer to the catalog. The injected sources were chosen to be empirical populations of artificial galaxies and stars taken from the deeper DES Deep Field observations (Hartley et al., 2021). Additionally, delta-function stars (delta functions convolved with a local PSF; Section 2.2.4 of Everett et al. (2022)) are injected, which we use for our synthetic stellar sample to avoid any potential galaxy contamination from incorrect Deep Field classifications. These synthetic objects are subject to the same systematic impacts from survey properties as physical objects would be. For this work we use all Balrog objects that pass our quality cuts. This consists of 7.4M total objects (90% Deep Field objects, 10% delta-stars). These objects were injected on a uniform grid on 2,041 randomly chosen tiles out of the 10,338 total Y3 tiles (for a coverage map see Figure 6 of Everett et al. 2022, and more information on the injected objects can be found in their Section 3).
@@ II.2 Balrog Synthetic Sources
III Methods
III.1 Overview and Notation
@@
Our algorithm uses SSI objects and survey property maps to derive a relation between the survey properties and the likelihood that a given object will be detected and classified correctly. This will allow us to take catalog level data ($C_{S}$ and $C_{G}$) and correct it based on the survey properties at a given on-sky location. Our correction will be based on the assumption that the synthetic objects will respond in the same way to survey property variations as real data. For Balrog objects, this assumption is appropriate based on Figure 11 of Everett et al. (2022). Since we know whether Balrog objects are truly stars or galaxies, we can use them to measure the rates at which objects are classified correctly. This measurement is critical for getting high accuracy counts of stars at faint magnitudes since misclassified galaxies are the dominant contributor to uncorrected faint star counts.
To start our correction, we therefore first use SSI objects to obtain the probability that a detected SSI star/galaxy will be classified correctly ($P\left(IC_{S}|IO_{S}\right)$ and $P\left(IC_{G}|IO_{G}\right)$). We assume that the probability that a detected star (galaxy) will be classified correctly is the same as that probability for an SSI star (galaxy):
@@
Classification probabilities and relative detection rates are used to obtain final corrected counts at each position ($FC_{S}$ & $FC_{G}$). We present two different types of corrections: one estimates what $C_{S}$ would have been with uniform survey properties, and one estimates what $O_{S}$ would have been with uniform survey properties. More details are given in Section III.3.
@@
III.2 Probability Calculations
Figure 2: This figure shows Balrog delta star correct classification rates before (blue) and after (orange) corrections as a function of the effective exposure time sum in the $i$-band. The green histogram shows the distribution of effective exposure time in the $i$-band across the DES footprint. Training was performed on an 80% subsample of the Balrog objects with the remaining sources used for testing. As well as magnitude cuts, a color cut to the Phoenix isochrone is made (see Section IV.1). A drop in standard deviation after corrections shows a mitigation of some of the variations caused by this survey property.
@@ III.2 Probability Calculations
As a specific example to illustrate our method, we present $P\left(IC_{S}|IO_{S}\right)$ in our faintest magnitude bin as a function of a single survey property in Figure 2. The green histogram shows the distribution of effective exposure time in the $i$-band for DES Y3. The blue (orange) line shows the ratio before (after) corrections of correctly classified SSI stars to all detected SSI stars as a function of the survey property value (Table 1). Since we only care about the variations in this rate, we divide each line by its average to center them at $1$. For corrections, this ratio of correctly classified SSI stars to all detected SSI stars is computed for all survey properties. We then take the survey property with the largest variation in the ratio and apply an empirical correction. This process is repeated iteratively until one of our termination criteria is met: either $\sigma<0.01$ for all survey properties or, to avoid overfitting and/or long runtimes, after 150 iterations (see Section V.3). To ensure that we are not just removing noise, we take 80% of our SSI stars to be a training sample and the remaining 20% to be a testing sample. All corrections were calculated on the training sample but then applied to the testing sample to make the plot. In this way, the drop in standard deviation from $\sigma=0.07$ to $\sigma=0.02$ indicates that our correction pipeline has reduced the impact variations in this survey property have on the stellar classification rate. We see $\sigma>0.01$ after corrections since the test set is independent from the training set.
@@
III.3 Algorithm Design
With classification probabilities and relative detection rates established, we now detail the correction algorithms introduced in Section III.1, which share identical steps except for their final stage.
The first step uses $P\left(C_{S}|O_{S}\right)$, $P\left(C_{G}|O_{G}\right)$, $C_{S}$, and $C_{G}$ to estimate the number of true stars (galaxies) that were detected, $RO_{S}$ ($RO_{G}$). To start, we note that the expected number of classified stars is:
$\langle C_{S}\rangle=O_{S}P\left(C_{S}|O_{S}\right)+O_{G}\left[P\left(C_{G}|O_{G}\right)-1\right],$ (2)
with an analogous equation for $\langle C_{G}\rangle$. For the above equation, we note that $\left[P\left(C_{G}|O_{G}\right)-1\right]=P\left(C_{S}|O_{G}\right)$.
These equations can be inverted to solve for the expected number of observed stars and observed galaxies ($RO_{S}$ and $RO_{G}$). This gives the maximum likelihood estimate for $RO_{S}$ and $RO_{G}$ of:
@@ III.3 Algorithm Design
IV Correction of the DES Data
In this section we present the results of applying our correction algorithm to DES data. To select a realistic dataset for this pipeline that will be applicable to a stream, we follow the study of the Phoenix stream by Tavangar et al. (2022). Phoenix is a $15^{\circ}$ long, $0.16^{\circ}$ wide, and dynamically cold stellar stream in the Southern Hemisphere (Balbinot et al., 2016; Shipp et al., 2018b; Tavangar et al., 2022). Phoenix shows density fluctuations on small scales, making it a well-suited candidate for studying potential perturbations. In addition to its intrinsic properties, the stream lies in the middle of the DES footprint, near a prominent survey depth feature at a right ascension (RA) of $\sim 30$ deg (Figure 1). This feature has increased depth and is one of the most readily apparent survey property features in the entire DES footprint. Therefore, we can use observations of the Phoenix stream to demonstrate the application of our pipeline and its effect on the generation of density maps.
@@ IV Correction of the DES Data
IV.1 Matched Filter Data selection
To derive corrections for a given stellar or galactic density map we want to place the same observational selection criteria on the injected objects as we would place on observations. In the case of the Phoenix stream, we use the matched filter from Tavangar et al. (2022). This filter is described in detail in their analysis, but generally it uses a synthetic isochrone to generate a selection region in color-magnitude space. The size of this region is defined by the expected width of the stellar population convolved with observational uncertainties. We use an isochrone from Bressan et al. (2012) as implemented in ugali (Bechtol et al., 2015; Drlica-Wagner et al., 2020) with an age of a =12.8=12.8 Gyr, metallcity of [Fe/H] =2.5=-2.5, and distance of d =17.4=17.4 kpc (Tavangar et al., 2022). In practice, extinction corrected data should be used to apply the isochrone cut and with apparent magnitudes then being used to apply corrections. This could be achieved by adding an extinction map as an additional survey property, but more Balrog objects would be needed to cover the full range of apparent magnitude space necessary for corrections, and binning would likely have to be done by color as well as magnitude. Since this paper is focused on proof of concept corrective power rather than legitimate DM constraints, we simplify our pipeline to only work with apparent magnitudes.
We also use a magnitude limit of $r<24.5$, where completeness falls to $\sim 20\%$ (see Everett et al. 2022 Figure 9). This can be compared to Shipp et al. (2018a) which use the same observations (DES Y3) but place a limit around one magnitude brighter ($g<23.5$). It is also $0.3$ mag fainter than the $r<24.2$ limit used in Tavangar et al. (2022) which uses DES Y6 data (roughly twice as many observations as DES Y3).
As mentioned in Section 2.1, in our correction algorithm we apply corrections to different magnitude bins. Our magnitude bins are as follows: $r\leq 22.9$, $22.9<r\leq 23.9$, $23.9<r\leq 24.5$. These magnitude bins were chosen as they had similar relative detection rate variations in testing sets when these variations were extrapolated out to expected values for a more current Balrog run. For more details, refer to Appendix E. These selections were placed in addition to the ones described in Section II. Throughout this section we show figures for these magnitude bins only, as this best illustrates the effectiveness of our method in extending the usable magnitude range for stream analyses; at brighter magnitudes the catalog is significantly more uniform.
IV.2 Algorithm Adjustments for Real Data